OpenStudy (anonymous):

Solve: log_{x}8 = 3/4

OpenStudy (anonymous):

\[\log_{x}8 = 3/4 \]

OpenStudy (anonymous):

8 = x^(3/4) 8^4 = x^3 8^(4/3) = x x = 16

OpenStudy (anonymous):

I think thats right... lol

OpenStudy (anonymous):

yes it is right!